# Token Distillation: Attention-Aware Input Embeddings for New Tokens

### v2

@@ B.9 Hyperparameters
B.10 Models
Appendix C Compute Budget & Runtime
@@
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Target 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 NTP (1 epoch) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 Token Distillation (1 epoch) 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 NTP (2 epoch) 61.0 66.1 65.4 65.9 68.7 56.7 61.5 58.9 63.0 Token Distillation (2 epoch) 63.2 68.5 68.9 67.7 71.8 58.4 63.6 61.3 65.4
@@
Comparing the computational budget with ZeTT (Minixhofer et al., 2024), ZeTT sees 3.2 billion tokens during hyper-network pretraining, while Token Distillation sees only 3.2 million tokens (for initializing 2,600 new tokens), less than 1% of ZeTT’s budget – still Token Distillation outperforms ZeTT. In return, ZeTT is faster at inference time (in our experiments on a single H100 80GB GPU, ZeTT took less than a minute). In future work, we believe that investigating a Token Distillation-style objective for hyper-network pretraining would be beneficial, as our experiments show that this significantly outperforms next-token prediction for learning new embeddings (ZeTT uses next-token prediction).
A note on downstream computational speedups.
Over-tokenization is not only associated with reduced downstream task performance (Rust et al., 2021; Ali et al., 2024) but also with increased computational cost due to longer sequence lengths (Ahia et al., 2023; Yamaguchi et al., 2024a). This computational cost is reduced when reducing over-tokenization by adding new (commonly occurring) tokens to the vocabulary. Importantly, for input embeddings, the reduction in computational cost is completely independent of the actual quality of the embeddings for the newly added tokens (for a given fixed sequence of text). The same computational speedups are achieved by using Random initialization and much better methods such as ZeTT (Minixhofer et al., 2024) or our proposed Token Distillation. Of course, the actual quality of any predictions made conditioned on a sequence containing such new tokens does heavily depend on the embedding quality. Therefore, in this work, we focus on direct evaluation of the quality of produced embeddings.
Appendix D Use of LLMs

### v3

@@ C.10 Hyperparameters
C.11 Continued Training Hyperparameters for Arabic Vocabulary Adaptation
For the continued-training setting in Table 5, we follow a lightweight vocabulary adaptation protocol based on AdaptiVocab (Nakash et al., 2025). We train the embedding matrices as well as the first and last Transformer layers. We use an effective batch size of 128 sequences of max length 2,048 and train for 1,000 steps (about 260M training tokens). We use AdamW with a learning rate of , 100 warmup steps, and cosine decay.
C.12 Models
Appendix D Compute Budget & Runtime
@@
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 NTP (1 epoch) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 Token Distillation (1 epoch) 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 NTP (2 epoch) 61.0 66.1 65.4 65.9 68.7 56.7 61.5 58.9 63.0 Token Distillation (2 epoch) 63.2 68.5 68.9 67.7 71.8 58.4 63.6 61.3 65.4
@@
Comparing the computational budget with ZeTT (Minixhofer et al., 2024), ZeTT sees 3.2 billion tokens during hyper-network pretraining, while Token Distillation sees only 3.2 million tokens (for initializing 2,600 new tokens), less than 1% of ZeTT’s budget – still Token Distillation outperforms ZeTT. In return, ZeTT is faster at inference time (in our experiments on a single H100 80GB GPU, ZeTT took less than a minute). In future work, we believe that investigating an Token Distillation-style objective for hyper-network pretraining would be beneficial, as our experiments show that this significantly outperforms next-token prediction for learning new embeddings (ZeTT uses next-token prediction).
Appendix E Use of LLMs